The car drove 8 miles in 12 minuets. what was the car's speed?
Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

The car drove 8 miles in 12 minuets. what was the car's speed?
---------
If you meant minutes:
d = r*t
r = d/t = 8/12
r = 2/3 mile/minute

Answer by solver91311(24713)



Hard to tell. How long is a "minuet"? I suppose it would depend on the actual tempo of the music, don't you? Besides learning how to spell, you might want to actually specify what units in which the speed should be expressed. Miles per hour? Miles per minute? Meters per second? Furlongs per hour? Nanometers per nanosecond?

In general however, rate is equal to distance divided by time. Stopping there with this problem would leave you with the rate in miles per minute (presuming you actually meant to write "12 minutes"). If the answer needs to be in some other unit of rate measurement, you need to apply the appropriate conversion factor.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism